OpenStudy (jetman24):

Which is the correct step below for using the grouping method to factor the following polynomial? 5xy + 15x - 2y - 6 A. x(5y + 15) - 1(2y + 6) B. 5x(y + 3) + 2(y + 3) C. 5x(y + 3) - 2(y + 3) D. 5x(y - 3) - 2(y - 3)

OpenStudy (jetman24):

medal to whoever solves this

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

(5xy + 15x) - (2y - 6) = 5x(y+3) -2(y+3) = (y+3)(5x-2) hopőe this will help you understanding easy

OpenStudy (anonymous):

yes sure than you check it these choice so choice C. is right,correct sure

OpenStudy (anonymous):

but i have made it for you step by step for you can understanding it easy right good luck bye bye

OpenStudy (anonymous):

the first 2 terms have 5x in common, so that has to come out; 5x(y+3)... The last two terms have a -2 in common, so that has to come out -2(y+3) and then just combine both parts of the equations: 5x(y+3)-2(y+3)..